find the following product
(x - 9)(x - 8)

Answer by stanbon(75887)

find the following product
(x - 9)(x - 8)
Use FOIL:
F: x*x = x^2
O: -8*x = -8x
I: -9*x = -9x
L: -9*-8 = 72
--------------
Combine those to get:
x^2-17x+72
=================
Cheers,
Stan H.